Circular No. 6417

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1995 O1 (HALE-BOPP) H. U. Käufl, European Southern Observatory (ESO), communicates:

“Comet C/1995 O1 was observed on June 5.25 UT with the ESO 3.6-m telescope (+ TIMMI, a focal reducer for the thermal infrared). Using the standard N filter (8–13 μm), the comet was clearly resolved and isophotes could be detected up to a radius of ∼ 5′′ from the nucleus. The images show a strong peak with a FWHM of 2′′.5–3′′ within a faint, diffuse cloud that is asymmetrically extended to the northeast. Using a photometric aperture of 7′′ × 7′′, the following fluxes have been provisionally derived: N, 6.85 Jy; N_{1} (7.9–8.9 μm), 1.46 Jy; N_{2} (9.14–10.34 μm), 2.82 Jy; N_{3} (11.8–13.0 μm), 6.74 Jy. The photometric results suggest an infrared dust color temperature of 200 K.”

M. S. Hanner, Jet Propulsion Laboratory, California Institute of Technology; and T. L. Hayward, Cornell University, report the detection with the 5-m Hale Telescope (+ SpectroCam-10) of strong silicate emission in the 8- to 13-μm spectrum of comet C/1995 O1 on June 11.3 UT. The feature is a factor of 1.8 above a 175-K continuum and shows an 11.25-μm peak from crystalline olivine similar to that seen in 1P/Halley. Nightly images of the inner coma during June 1–12 reveal a prominent jet extending 5′′ north of the nucleus on June 4.5 and 9.5, while the jet was considerably weaker or absent on all other nights. On June 4.5, the total flux in a 10′′ circle was 5 Jy at 10.3 μm.

Visual m_{1} and coma-diameter estimates (cf. IAUC 6408): May 28.17 UT, 6.8, 10′ × 14′ (J. M. Trigo, Castellon, Spain, 7×50 binoculars); June 6.44, 6.3, – (D. Seargent, The Entrance, N.S.W., 10×50 binoculars); 8.37, 6.6, 6′.5 (C. E. Spratt, Victoria, BC, 0.10-m refractor); 10.96, 6.1, 14′ (M. V. Zanotta, San Marco Pass, Italy, 7×42 binoculars); 12.11, 6.3, 14′ (R. Lourencon, Jundiai, São Paulo, Brazil, 20×80 binoculars); 12.99, 6.4, 9′ (R. J. Bouma, Groningen, The Netherlands, 7×50 binoculars).

COMET 65P/GUNN Total visual magnitude estimates (cf. IAUC 6392): May 13.28 UT, 12.7

(A. Hale, Cloudcroft, NM, 0.41-m reflector); 19.01, 13.6 (K. Sárneczky, Ráktanya, Hungary, 0.44-m reflector); 25.02, 12.3 (J. Carvajal, Teruel, Spain, 0.44-m reflector); June 8.98, 12.0 (K. Hornoch, Lelekovice, Czech Republic, 0.35-m reflector).

1996 June 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green